# Safety, Relative Tightness and the Probabilistic Frame Rule

### v6

@@ 2 Mathematical preliminaries
A (discrete probability) distribution on a set $A$ is a function $d\colon A\to[0,1]$ that satisfies
@@
The support $\mathsf{Supp}(d)$ is necessarily a countable set. We write $\mathcal{D}(A)$ for the set of all distributions on $A$. A function $f:A\to B$ induces a function $f_{!}:\mathcal{D}(A)\to\mathcal{D}(B)$ that maps every $d\in\mathcal{D}(A)$ to its pushforward
@@
For an arbitrary set $A$, an $A$-valued random variable is a function $\Sigma:\Omega\to\mathsf{State}$, where $\Omega$ is a (discrete) sample space, namely a (necessarily countable) set equipped with a probability distribution $p_{\Omega}\colon\Omega\to[0,1]$, for which we require the additional (not strictly necessary, but natural) condition that $\mathsf{Supp}(d)=\Omega$; that is, there are no zero-probability sample-space elements. It will sometimes be useful to change sample space along a morphism of sample spaces, which is a function $q:\Omega^{\prime}\to\Omega$ satisfying, $q_{!}(p_{\Omega^{\prime}})=p_{\Omega}$. Since sample spaces have no zero-probability elements, every morphism of sample spaces is surjective. To emphasise this point in the sequel, we shall write $q:\Omega^{\prime}\twoheadrightarrow\Omega$ for sample space morphisms.
Given a random variable $S:\Omega\to A$, we write $p_{S}:A\to[0,1]$ for the distribution of $S$, given by the pushforward probability distribution $S_{!}(p_{\Omega})$ on $A$. We define the support of a random variable $S$ by $\mathsf{Supp}(S):=\mathsf{Supp}(p_{S})$.
@@ 4 Assertions
Since the values of variables are in general randomly distributed, we interpret the above equality as stating that $X$ and $Y$ take the same value with probability $1$; that is, considered as random variables, $X$ and $Y$ are almost surely equal.
@@ 6 The frame rule
Since both $C,\sigma$ and $C,{\sigma\!\!\restriction_{\mathrm{FV}(\Phi)}}$ are terminating, they induce probability distributions $p_{C,\sigma}$ and $p_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$ on their terminal states. Our next goal is to relate these two distributions. By Lemma 6.6, every terminal path $\pi\in\mathcal{TP}_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$, induces a corresponding terminal path $\langle\sigma\rangle\pi\in\mathcal{TP}_{C,\sigma}$ with $\mathsf{Pr}(\langle\sigma\rangle\pi)=\mathsf{Pr}(\pi)$. Since $C,{\sigma\!\!\restriction_{\mathrm{FV}(\Phi)}}$ is terminating, the probabilities $\mathsf{Pr}(\pi)$ of all $\pi\in\mathcal{TP}_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$ add up to $1$, hence so do all probabilities $\mathsf{Pr}(\langle\sigma\rangle\pi)$. Thus every terminal path in $\mathcal{TP}_{C,\sigma}$ is necessarily of the form $\langle\sigma\rangle\pi$ for some $\pi\in\mathcal{TP}_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$. It follows that the induced probabilities on terminal states, $p_{C,\sigma}$ and $p_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$, are related to each other by the pushforward property
@@
where $\top$ is the true assertion. Recall that $X=0\vee X=1]$ says that the stated property holds with probability $1$. Under our definition of partial correctness (Definition 5.1), this specification is not partially correct because it fails the safety guarantee: the empty random state satisfies the precondition, but execution from the empty state faults. If, however, the safety guarantee is dropped from the definition of partial correctness, then the above specification becomes correct. In any random state $\Sigma$ from which the program $X{:=}X\ \mathsf{mod}\ 2$ terminates, the postcondition $\{[X=0\vee X=1]\}$ is indeed true.

### v7

@@ 2 Mathematical preliminaries
A (discrete probability) distribution on a set AA is a function $d\colon A\to[0,1]$ that satisfies
@@
The support $\mathsf{Supp}(d)$ is necessarily a countable set. We write $\mathcal{D}(A)$ for the set of all distributions on AA. A function $f:A\to B$ induces a function $f_{!}:\mathcal{D}(A)\to\mathcal{D}(B)$ that maps every $d\in\mathcal{D}(A)$ to its pushforward
@@
For an arbitrary set AA, an AA-valued random variable is a function $\Sigma:\Omega\to\mathsf{State}$, where $\Omega$ is a (discrete) sample space, namely a (necessarily countable) set equipped with a probability distribution $p_{\Omega}\colon\Omega\to[0,1]$, for which we require the additional (not strictly necessary, but natural) condition that $\mathsf{Supp}(d)=\Omega$; that is, there are no zero-probability sample-space elements. It will sometimes be useful to change sample space along a morphism of sample spaces, which is a function $q:\Omega^{\prime}\to\Omega$ satisfying, $q_{!}(p_{\Omega^{\prime}})=p_{\Omega}$. Since sample spaces have no zero-probability elements, every morphism of sample spaces is surjective. To emphasise this point in the sequel, we shall write $q:\Omega^{\prime}\twoheadrightarrow\Omega$ for sample space morphisms.
Given a random variable $S:\Omega\to A$, we write $p_{S}:A\to[0,1]$ for the distribution of $S$, given by the pushforward probability distribution $S_{!}(p_{\Omega})$ on AA. We define the support of a random variable $S$ by $\mathsf{Supp}(S):=\mathsf{Supp}(p_{S})$.
@@ 4 Assertions
Since the values of variables are in general randomly distributed, we interpret the above equality as stating that $X$ and $Y$ take the same value with probability 11; that is, considered as random variables, $X$ and $Y$ are almost surely equal.
@@ 6 The frame rule
Since both $C,\sigma$ and $C,{\sigma\!\!\restriction_{\mathrm{FV}(\Phi)}}$ are terminating, they induce probability distributions $p_{C,\sigma}$ and $p_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$ on their terminal states. Our next goal is to relate these two distributions. By Lemma 6.6, every terminal path $\pi\in\mathcal{TP}_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$, induces a corresponding terminal path $\langle\sigma\rangle\pi\in\mathcal{TP}_{C,\sigma}$ with $\mathsf{Pr}(\langle\sigma\rangle\pi)=\mathsf{Pr}(\pi)$. Since $C,{\sigma\!\!\restriction_{\mathrm{FV}(\Phi)}}$ is terminating, the probabilities $\mathsf{Pr}(\pi)$ of all $\pi\in\mathcal{TP}_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$ add up to 11, hence so do all probabilities $\mathsf{Pr}(\langle\sigma\rangle\pi)$. Thus every terminal path in $\mathcal{TP}_{C,\sigma}$ is necessarily of the form $\langle\sigma\rangle\pi$ for some $\pi\in\mathcal{TP}_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$. It follows that the induced probabilities on terminal states, $p_{C,\sigma}$ and $p_{C,\,{\sigma\!\restriction_{\mathrm{FV}(\Phi)}}}$, are related to each other by the pushforward property
@@
where $\top$ is the true assertion. Recall that $X=0\vee X=1]$ says that the stated property holds with probability 11. Under our definition of partial correctness (Definition 5.1), this specification is not partially correct because it fails the safety guarantee: the empty random state satisfies the precondition, but execution from the empty state faults. If, however, the safety guarantee is dropped from the definition of partial correctness, then the above specification becomes correct. In any random state $\Sigma$ from which the program $X{:=}X\ \mathsf{mod}\ 2$ terminates, the postcondition $\{[X=0\vee X=1]\}$ is indeed true.